OpenStudy (anonymous):

If x+1/x=2, then find the principal value of sin-1x.

OpenStudy (anonymous):

hmm?

OpenStudy (anonymous):

$$x+\frac1x=2\\x^2+1=2x\\x^2-2x+1=0\\(x-1)^2=0\\x=1$$Can you figure out \(\arcsin 1\)?

OpenStudy (anonymous):